# A stochastic preconditioned Douglas-Rachford splitting method for saddle-point problems

### v2

@@ 3 Almost sure convergence of stochastic PDR
Inserting $u^{*}=(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})$ a fixed point of $\mathcal{T}$ in (3.5). By Lemma 2, $(x^{*},y^{*})$ is a saddle point of (1.1). Then it follows for every $(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})\in\mathbb{F}$, $\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),(x^{*},y^{*}))\geq 0$. Denoting
$\Delta^{k}=\|u^{k}-u^{*}\|^{2}_{\mathbf{M}_{p}},$
and employing Lemma 4 as well as $\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{p}}=\|\mathcal{T}u^{k}-u^{k}\|^{2}_{\mathbf{M}}$, we have
$\mathbb{E}(\Delta^{k+1}|\mathfrak{F}_{k})=\mathbb{E}^{k}\Delta^{k+1}\leq\Delta^{k}-\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{p}}=\Delta^{k}-\|\mathcal{T}u^{k}-u^{k}\|^{2}_{\mathbf{M}}.$ (3.10)
Using the Robbins–Siegmund theorem [21] (see also [11, 1]) on (3.10), we conclude that
$\sum_{k=0}^{\infty}\|\mathcal{T}u^{k}-u^{k}\|^{2}_{\mathbf{M}}<\infty,\quad\|\mathcal{T}u^{k}-u^{k}\|^{2}_{\mathbf{M}}\rightarrow 0,$ (3.11)
almost surely in $\Omega$. Recalling the definition of $\mathbf{M}_{p}$, we immediately see that also $\|\mathcal{T}u^{k}-u^{k}\|_{\mathbf{M}_{p}}\to 0$ as $k\to\infty$ almost surely. Using further that $\|u^{k}-u^{*}\|_{\mathbf{M}_{p}}=\|\mathbf{M}_{p}^{1/2}(u^{k}-u^{*})\|$ for each $k$ and applying Proposition 2.3 in [11], we conclude that $\{\mathbf{M}_{p}^{1/2}u^{k}\}_{k}$ is bounded almost surely and that the limit
@@
exists almost surely. In the following, we will establish the almost sure weak convergence of $\{u^{k}\}_{k}$. Our first goal is to show that $\{\tilde{u}^{k}\}_{k}$ converges weakly to a fixed point of $\mathcal{T}$. We will begin with proving the almost sure weak convergence of $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ with Opial’s lemma [2] (Lemma 21). This can be done by showing that almost surely, $\|\mathbf{M}_{p}^{1/2}(\tilde{u}^{k+1}-u^{*})\|$ converges and every weak sequential limit point of $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k+1}\}_{k}$ belongs to $\mathbf{M}_{p}^{1/2}\mathbb{F}$. According to the triangle inequality, we have
$\big{|}\|\mathbf{M}_{p}^{1/2}(\tilde{u}^{k+1}-u^{*})\|-\|\mathbf{M}_{p}^{1/2}(u^{k}-u^{*})\|\big{|}\leq\|\mathbf{M}_{p}^{1/2}(\tilde{u}^{k+1}-u^{k})\|=\|\mathcal{T}u^{k}-u^{k}\|_{\mathbf{M}_{p}},$
so
@@
almost surely as $k\to\infty$. In particular, $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ is bounded almost surely. Next, we show that for each event such that (3.11), (3.12) and (3.13) hold, each weak accumulation point of $\{\tilde{u}^{k}\}_{k}$ is a fixed point. For that purpose, note that due to (3.12), we have for the sequence $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ and each $\mathbf{M}_{p}^{1/2}u^{*}\in\mathbf{M}_{p}^{1/2}\mathbb{F}$ that $\|\mathbf{M}_{p}^{1/2}\tilde{u}^{k}-\mathbf{M}_{p}^{1/2}u^{*}\|$ converges. Now, let $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k_{j}}\}_{j}$ be a weakly converging subsequence. Then, as $\|\mathbf{M}_{p}^{1/2}(\mathcal{T}u^{k}-u^{k})\|\to 0$ as $k\to\infty$, $\{\mathbf{M}_{p}^{1/2}u^{k_{j}-1}\}_{j}$ converges weakly to the same limit and looking at the definition of $\mathbf{M}_{p}$, we conclude that $\bigl{\{}\bigl{(}(N_{1}-I)x^{k_{j}-1},(N_{2}-I)y^{k_{j}-1},\bar{x}^{k_{j}-1},\bar{y}^{k_{j}-1}\bigr{)}\bigr{\}}_{j}$ converges weakly. We further have that $\bigl{(}(N_{1}-I)(\tilde{x}^{k_{j}}-x^{k_{j}-1}),(N_{2}-I)(\tilde{y}^{k_{j}}-y^{k_{j}-1}),\tilde{\bar{x}}^{k_{j}}-\bar{x}^{k_{j}-1},\tilde{\bar{y}}^{k_{j}}-\bar{y}^{k_{j}-1}\bigr{)}\to 0$ as $j\to\infty$ such that with Lemma 2 (ii), we obtain that $\{(\tilde{x}^{k_{j}},\tilde{y}^{k_{j}},\tilde{\bar{x}}^{k_{j}},\tilde{\bar{y}}^{k_{j}})\}$ converges weakly to a fixed point $(x^{**},y^{**},\bar{x}^{**},\bar{y}^{**})$ of $\mathcal{T}$. In particular, by Lemma 2 (i), $(x^{**},y^{**})$ is a saddle point for (1.1) and $x^{**}+\sigma\mathcal{K}^{*}y^{**}=\bar{x}^{**},y^{**}-\tau\mathcal{K}x^{**}=\bar{y}^{**}$. It is then immediate that the weak limit of $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k_{j}}\}_{j}$ belongs to $\mathbf{M}_{p}^{1/2}\mathbb{F}$. By Opial’s lemma, we have that the whole sequence $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ converges weakly to an element in $\mathbf{M}_{p}^{1/2}\mathbb{F}$. Repeating the above arguments with $k_{j}=j+1$, it follows that $\{\tilde{u}^{k}\}_{k}$ converges weakly to the fixed point $u^{**}$ of $\mathcal{T}$. Finally, let us turn to the almost sure weak convergence of $\{u^{k}\}_{k}$. First, note that since for fixed $i=1,\ldots,n$, the probability $p_{i}$ of $i_{k}=i$ is positive and all $i_{k}$ are independent random variables, the probability of $i_{k}=i$ only occurring finitely many times has to be zero. For that reason, almost surely, for each $i=1,\ldots,n$, the sequences $\{k_{i,j}\}_{j}$ such that $k_{i,j}=\max\ \{k:k\leq j,\ i_{k-1}=i\}$, i.e., the sequences that yield the index of the last activation of $i$, are increasing and satisfy $\lim_{j\to\infty}k_{i,j}=\infty$. Fix an event for which $\{\tilde{u}^{k}\}_{k}$ converges weakly to a fixed point $u^{**}=(x^{**},y^{**},\bar{x}^{**},\bar{y}^{**})$ of $\mathcal{T}$ and for which $k_{i,j}$ tends to infinity as $j\to\infty$ for all $i=1,\ldots,n$. We will show that $\{u^{k}\}_{k}$ converges weakly, as $k\to\infty$, to $u^{*}$. For that purpose, observe that by (3.2), we have $x^{k}=\tilde{x}^{k}\to x^{**}$, $x^{k}_{\operatorname{test}}=\tilde{x}^{k}+\tilde{\bar{x}}^{k}-\bar{x}^{k-1}\to x^{**}$ and $\bar{x}^{k}=\tilde{\bar{x}}^{k}\to\bar{x}^{**}$ in the weak sense as $k\to\infty$. By (3.3) and the definition of $k_{i,j}$, we see that $y^{j}_{i}=\tilde{y}^{k_{i,j}}_{i}$, $y^{j}_{\operatorname{test},i}=\tilde{y}^{k_{i,j}}_{i}+\tilde{\bar{y}}^{k_{i,j}}_{i}-\bar{y}^{(k_{i,j})-1}_{i}$ and $\bar{y}^{j}_{i}=\tilde{\bar{y}}^{k_{i,j}}_{i}$ for each $j$. Hence, letting $j\to\infty$, we see that $y^{j}_{i}\to y^{**}_{i}$, $y^{j}_{\operatorname{test},i}\to y^{**}_{i}$ and $\bar{y}^{j}_{i}\to\bar{y}^{**}_{i}$ in the weak sense. As the events for which this follows occur almost surely, the proof is complete. ∎
@@
$N_{1}>I;N_{2}>I$ Easy to prove with [1] Presented in this paper
$N_{1}\geq I,N_{1}\neq I;N_{2}=I$ Presented in this paper Presented in this paper
As shown in Table 1, our framework covers some results in existing research. The commonly used preconditioners for stochastic PDR are the case $N_{1}\geq I,N_{1}\neq I$ and $N_{2}=I$, leading to difficulties in the proof. Next, let us turn to the convergence rate of the restricted primal-dual gap, which is also a challenge due to the preconditioners and the transitional variables.
4 Sublinear convergence of stochastic PDR for restricted primal-dual gap
@@
$\mathbb{E}[\Delta^{k}]\leq\Delta^{0}=\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p}},\quad\sum_{i=0}^{k}\mathbb{E}[\|u^{i+1}-u^{i}\|^{2}_{\mathbf{M}_{p}}]\leq\Delta^{0}.$ (4.1)
Let us begin with the following lemma which gives the control of $\mathcal{T}u^{k}$ by the weighted $\mathbf{M}_{p}$ norm of $u^{k}$. Applying Lemma 2.5 in [8] with $\mathbf{M}_{p}$, we have the following estimate.
@@
where the constant $C$ is only related to $N_{1}$, $N_{2}$, $\sigma$, $\tau$ and $\mathcal{K}$.
The following simple lemma can deal with the main challenge in $y^{k+1}_{\operatorname{test}}$ for the gap estimate.
@@
    $\mathbb{E}\varphi(s_{i}^{k+1})=\sum_{n=0}^{k}\mathbb{E}[p_{i}(1-p_{i})^{n}\varphi(\tilde{s}_{i}^{k+1-n})]+(1-p_{i})^{k+1}\varphi(s_{i}^{0}),$
@@
    $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-y^{*}\|^{2}\leq n(3C+3\tau)\Delta^{0}+\|y^{0}_{\operatorname{test}}-y^{*}\|^{2},$ where $C$ is as in Lemma 5.
@@
    $\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}\leq D\Delta^{0}$, where $D>0$ is a positive constant.

### v3

@@ 3 Almost sure convergence of stochastic RPDR
Inserting $u^{*}=(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})$ a fixed point of $\mathcal{T}_{R}$ in (3.6). By Lemma 2, $(x^{*},y^{*})$ is a saddle point of (1.1). Then it follows for every $(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})\in\mathbb{F}$, $\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),(x^{*},y^{*}))\geq 0$. Employing Lemma 4 as well as $\mathbf{M}_{p,\varrho_{k+1}}\preceq\mathbf{M}_{p,\varrho_{k}}$ we have
$\mathbb{E}^{k}\|u^{k+1}-u^{*}\|^{2}_{\mathbf{M}_{p,\varrho_{k+1}}}$ $\leq\mathbb{E}^{k}\|u^{k+1}-u^{*}\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}$
$\leq\|u^{k}-u^{*}\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}-\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}_{p}},$
denoting $\Delta^{k}=\|u^{k}-u^{*}\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}$, and due to $\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}_{p}}=\|\mathcal{T}_{R}u^{k}-u^{k}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}}$, we get,
$\mathbb{E}(\Delta^{k+1}|\mathfrak{F}_{k})=\mathbb{E}^{k}\Delta^{k+1}$ $\leq\Delta^{k}-\|\mathcal{T}_{R}u^{k}-u^{k}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}}.$ (3.11)
Using the Robbins–Siegmund theorem [38] (see also [15, 1]) on (3.11), we conclude that $\sum_{k=0}^{\infty}\|\mathcal{T}_{R}u^{k}-u^{k}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}}<\infty$. Due to $\rho_{k},\rho_{k,x},\rho_{k,y}$ are non-decreasing, thus $\frac{2-\rho_{k}}{\rho_{k}^{2}}$, $\frac{2-\rho_{k,x}}{\rho_{k,x}^{2}}$ and $\frac{2-\rho_{k,y}}{\rho_{k,y}^{2}}$ are decreasing and the lower bound is $\frac{2-\rho_{u}}{\rho_{u}^{2}}$. Then it follows that
$\sum_{k=0}^{\infty}\frac{2-\rho_{u}}{\rho_{u}^{2}}\|\mathcal{T}_{R}u^{k}-u^{k}\|^{2}_{\mathbf{M}}<\infty,\quad\|\mathcal{T}_{R}u^{k}-u^{k}\|^{2}_{\mathbf{M}}\rightarrow 0,$ (3.12)
almost surely in $\Omega$. Recalling the definition of $\mathbf{M}_{p}$, we immediately see that also $\|\mathcal{T}_{R}u^{k}-u^{k}\|_{\mathbf{M}_{p}}\to 0$ as $k\to\infty$ almost surely. Still with the Robbins–Siegmund theorem or Proposition 2.3 in [15], we conclude that $\Delta^{k}=\|u^{k}-u^{*}\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}$ converges to a finite random variable almost surely. Moreover, $\|u^{k}-u^{*}\|_{\mathbf{M}_{p}}$ also converges almost surely by the non-decreasing and convergence of $\rho_{k},\rho_{k,x},\rho_{k,y}$. Using further that $\|u^{k}-u^{*}\|_{\mathbf{M}_{p}}=\|\mathbf{M}_{p}^{1/2}(u^{k}-u^{*})\|$ for each $k$, we conclude that $\{\mathbf{M}_{p}^{1/2}u^{k}\}_{k}$ is also bounded almost surely and that the limit
@@
exists almost surely. The existence and boudnedness of $\mathbf{M}_{p}^{1/2}$ can be guaranteed by the positive semi-definiteness of $N_{1}-I$ and $N_{2}-I$ and the positive definiteness of $\mathbf{P}$. In the following, we will establish the almost sure weak convergence of $\{u^{k}\}_{k}$.
Our first goal is to show that $\{\tilde{u}^{k}\}_{k}$ converges weakly to an $\mathbb{F}$-valued random variable, i.e., a fixed point of $\mathcal{T}_{R}$. We will begin with proving the almost sure weak convergence of $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ with Opial’s lemma [3, Lemma 21]. This can be done by showing that almost surely, $\|\mathbf{M}_{p}^{1/2}(\tilde{u}^{k+1}-u^{*})\|$ converges and every weak sequential limit point of $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k+1}\}_{k}$ is a $(\mathbf{M}_{p}^{1/2}\mathbb{F})$-valued random variable. Here $\mathbf{M}_{p}^{1/2}\mathbb{F}$ is defined as follows
$\mathbf{M}_{p}^{1/2}\mathbb{F}:=\{((N_{1}-I)^{\frac{1}{2}}x^{*}/\sigma^{\frac{1}{2}},(N_{2}-I)^{\frac{1}{2}}\mathbf{P}^{-\frac{1}{2}}y^{*}/\tau^{\frac{1}{2}},\bar{x}^{*}/\sigma^{\frac{1}{2}},\mathbf{P}^{-\frac{1}{2}}\bar{y}^{*}/\tau^{\frac{1}{2}})\ |\ u^{*}=(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})\in\mathbb{F}\},$
with $\mathbf{P}^{-\frac{1}{2}}=\operatorname{Diag}[p_{1}^{-\frac{1}{2}}I,\ldots,p_{n}^{-\frac{1}{2}}I]$. According to the triangle inequality, (3.12), we have
$\big{|}\|\mathbf{M}_{p}^{1/2}(\tilde{u}^{k+1}-u^{*})\|-\|\mathbf{M}_{p}^{1/2}(u^{k}-u^{*})\|\big{|}\leq\|\mathbf{M}_{p}^{1/2}(\tilde{u}^{k+1}-u^{k})\|=\|\mathcal{T}_{R}u^{k}-u^{k}\|_{\mathbf{M}_{p}}\rightarrow 0.$
Combining the fact of (3.13), we derive,
@@
almost surely as $k\to\infty$. In particular, $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ is bounded almost surely. Next, we show that for each event such that (3.12), (3.13) and (3.14) hold, each weak accumulation point of $\{\tilde{u}^{k}\}_{k}$ is a fixed point. For that purpose, note that (3.14), we have for the sequence $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ and each $\mathbf{M}_{p}^{1/2}u^{*}$ being a $(\mathbf{M}_{p}^{1/2}\mathbb{F})$-valued random variable that $\|\mathbf{M}_{p}^{1/2}\tilde{u}^{k}-\mathbf{M}_{p}^{1/2}u^{*}\|$ converges. Now, let $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k_{j}}\}_{j}$ be a weakly converging subsequence. Then, as $\|\mathbf{M}_{p}^{1/2}(\mathcal{T}_{R}u^{k}-u^{k})\|\to 0$ as $k\to\infty$, $\{\mathbf{M}_{p}^{1/2}u^{k_{j}-1}\}_{j}$ converges weakly to the same limit and looking at the definition of $\mathbf{M}_{p}$, we conclude that $\bigl{\{}\bigl{(}(N_{1}-I)x^{k_{j}-1},(N_{2}-I)y^{k_{j}-1},\bar{x}^{k_{j}-1},\bar{y}^{k_{j}-1}\bigr{)}\bigr{\}}_{j}$ converges weakly. We further have that $\bigl{(}(N_{1}-I)(\tilde{x}^{k_{j}}-x^{k_{j}-1}),(N_{2}-I)(\tilde{y}^{k_{j}}-y^{k_{j}-1}),\tilde{\bar{x}}^{k_{j}}-\bar{x}^{k_{j}-1},\tilde{\bar{y}}^{k_{j}}-\bar{y}^{k_{j}-1}\bigr{)}\to 0$ as $j\to\infty$ such that with Lemma 2 (ii), we obtain that $\{(\tilde{x}^{k_{j}},\tilde{y}^{k_{j}},\tilde{\bar{x}}^{k_{j}},\tilde{\bar{y}}^{k_{j}})\}$ converges weakly to a fixed point $(x^{**},y^{**},\bar{x}^{**},\bar{y}^{**})$ of $\mathcal{T}_{R}$. In particular, by Lemma 2 (i), $(x^{**},y^{**})$ is a saddle point for (1.1) and $x^{**}+\sigma\mathcal{K}^{*}y^{**}=\bar{x}^{**},y^{**}-\tau\mathcal{K}x^{**}=\bar{y}^{**}$. It is then immediate that the weak limit of $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k_{j}}\}_{j}$ is a $(\mathbf{M}_{p}^{1/2}\mathbb{F})$-valued random variable. By Opial’s lemma, we have that the whole sequence $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ converges weakly to a $(\mathbf{M}_{p}^{1/2}\mathbb{F})$-valued random variable, implying that $(\tilde{\bar{x}}^{k},\tilde{\bar{y}}^{k})$ converges weakly to $(\bar{x}^{**},\bar{y}^{**})$. Then by (3.12) again, we have $\bigl{(}(N_{1}-I)(\tilde{x}^{k}-x^{k-1}),(N_{2}-I)(\tilde{y}^{k}-y^{k-1}),\tilde{\bar{x}}^{k}-\bar{x}^{k-1},\tilde{\bar{y}}^{k}-\bar{y}^{k-1}\bigr{)}\to 0$ as $k\to\infty$. Applying Lemma 2 (ii) again, we conclude that $\{\tilde{u}^{k}\}_{k}$ converges weakly to a $\mathbb{F}$-valued random variable.
Finally, let us turn to the almost sure weak convergence of $\{u^{k}\}_{k}$. First, note that since for fixed $i=1,\ldots,n$, the probability $p_{i}$ of $i_{k}=i$ is positive and all $i_{k}$ are independent random variables, the probability of $i_{k}=i$ only occurring finitely many times has to be zero. For that reason, almost surely, for each $i=1,\ldots,n$, the sequences $\{k_{i,j}\}_{j}$ such that $k_{i,j}=\max\ \{k:k\leq j,\ i_{k-1}=i\}$, i.e., the sequences that yield the index of the last activation of $i$, are increasing and satisfy $\lim_{j\to\infty}k_{i,j}=\infty$. Fix an event for which $\{\tilde{u}^{k}\}_{k}$ converges weakly to a fixed point $u^{**}=(x^{**},y^{**},\bar{x}^{**},\bar{y}^{**})$ of $\mathcal{T}_{R}$ and for which $k_{i,j}$ tends to infinity as $j\to\infty$ for all $i=1,\ldots,n$. We will show that $\{u^{k}\}_{k}$ converges weakly, as $k\to\infty$, to $u^{*}$. For that purpose, observe that by (3.3), we have $x^{k}=\tilde{x}^{k}\to x^{**}$, $x^{k}_{\operatorname{test}}=\tilde{x}^{k}+\tilde{\bar{x}}^{k}-\bar{x}^{k-1}\to x^{**}$ and $\bar{x}^{k}=\tilde{\bar{x}}^{k}\to\bar{x}^{**}$ in the weak sense as $k\to\infty$. By (3.4) and the definition of $k_{i,j}$, we see that $y^{j}_{i}=\tilde{y}^{k_{i,j}}_{i}$, $y^{j}_{\operatorname{test},i}=\tilde{y}^{k_{i,j}}_{i}+\tilde{\bar{y}}^{k_{i,j}}_{i}-\bar{y}^{(k_{i,j})-1}_{i}$ and $\bar{y}^{j}_{i}=\tilde{\bar{y}}^{k_{i,j}}_{i}$ for each $j$. Hence, letting $j\to\infty$, we see that $y^{j}_{i}\to y^{**}_{i}$, $y^{j}_{\operatorname{test},i}\to y^{**}_{i}$ and $\bar{y}^{j}_{i}\to\bar{y}^{**}_{i}$ in the weak sense. Since these events occur almost surely, the proof is complete. ∎
@@
$N_{1}\succ I;N_{2}\succ I$ Easy to prove with [1] Presented in this paper
$N_{1}\succeq I,N_{1}\neq I;N_{2}=I$ Presented in this paper Presented in this paper
As shown in Table 1, our framework covers some results in existing research. The commonly used preconditioners for stochastic (and relaxed) PDR are the cases of $N_{1}\succeq I,N_{1}\neq I$ and $N_{2}=I$, which brings out the degeneracy of $N_{1}-I$ and leads to the difficulties in this proof. For instance, the symmetric Gauss-Seidel preconditioner is widely used for an elliptic operator $T$ as in (2.1) and $M-T=N_{1}-I$ is only positive semidefinite [10, Proposition 2.13]. Next, let us turn to the convergence rate of the restricted primal-dual gap with respect to the transitional variables $(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$. It is also a challenge due to the preconditioners and particularly because of the transitional variables, which are the main differences compared to the gap estimate reported in [1]. Since the relations derived from (3.11) are concerning $u^{k}$, i.e., $(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})$, without any relations of transitional variables $(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$, we introduce Lemma 5, Lemma 6 and Lemma 7 to address this issue.
4 Sublinear convergence of stochastic RPDR for restricted primal-dual gap
@@
$\mathbb{E}[\Delta^{k}]\leq\Delta^{0}=\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p,\varrho_{0}}},\ \ \sum_{i=0}^{k}\frac{2-\rho_{u}}{\rho_{u}^{2}}\mathbb{E}[\|u^{i+1}-u^{i}\|^{2}_{\mathbf{M}_{p}}]\leq\sum_{i=0}^{k}\mathbb{E}[\|u^{i+1}-u^{i}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}_{p}}]\leq\Delta^{0}.$ (4.1)
Since SRPDR involves $u^{k+1}_{t}=\mathcal{T}u^{k}$ in the iteration, let us begin with the following lemma which gives the control of $u^{k+1}_{t}$ by the weighted $\mathbf{M}_{p}$ norm of $u^{k}$, where $u^{k}$ is the iteration sequence of SRPDR. Applying [10, Lemma 2.5] with $\mathbf{M}_{p}$ norm, we have the following estimate.
@@
where $\mathcal{T}$ is the iteration operator of PDR, the constant $C$ is only related to $N_{1}$, $N_{2}$, $\sigma$, $\tau$ and $\mathcal{K}$.
The following two lemmas can address the main challenges in $y^{k+1}_{\operatorname{test}}$ for the gap estimate.
@@
    $\mathbb{E}\varphi(s_{i}^{k+1})=\mathbb{E}[\sum_{j=0}^{k}p_{i}(1-p_{i})^{j}\varphi(\tilde{s}_{i}^{k+1-j})]+(1-p_{i})^{k+1}\varphi(s_{i}^{0}),$
@@
    $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-y^{*}\|^{2}\leq\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|y^{0}_{\operatorname{test}}-y^{*}\|^{2},$ where $C$ is as in Lemma 5.
@@
    $\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}\leq\frac{D\rho^{2}_{u}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}$, where $D>0$ is a positive constant.